OpenStudy (anonymous):

Solve the system 2x + 3y = 3 and 3x – 2y = 11 by using graph paper or graphing technology. What is the solution to the system? (–3, 3) (–1, –7) (1, –4) (3, –1)

OpenStudy (anonymous):

Which are you using, paper or calculator?

OpenStudy (anonymous):

paper..

OpenStudy (anonymous):

Ok, then: 1) solve both systems for y in terms of x (put x's on one side, y's on the other; divide by the coefficient of y 2) Plot both lines 3) The solution is the point at which the two lines intersect

OpenStudy (anonymous):

If you need elaboration on any step, just let me know.

OpenStudy (anonymous):

2*(2x+3y)+3*(3x-2y)=6+33 13x=39 x=3 thus y=-1

OpenStudy (anonymous):

@chanh_chung That is neither using a graph, nor using a graphing technology. Please read questions before responding.

OpenStudy (anonymous):

@andreadesirepen Did you get the equations yet?

OpenStudy (anonymous):

yes.. thankyou so much

OpenStudy (anonymous):

Not a problem.

OpenStudy (anonymous):

Not a problem.